OpenStudy (anonymous):

Let f(x) = 5/b+5 - 2/b . Find the domain and simplify.

OpenStudy (anonymous):

you cannot have \[f(x)\] on one side and an expression in b on the other. is this \[f(x)=\frac{5}{x+5}-\frac{2}{x}\]?

OpenStudy (anonymous):

if so the implied domain is all numbers except \[x=-5,x=0\] because these would make the denominator 0,and you cannot divide by 0

OpenStudy (anonymous):

it has a b on the other side \[f(x) = \frac{5}{b+5} - \frac{2}{b}\]

OpenStudy (anonymous):

then it makes no sense. this is the second time i saw this exact same problem. i am wondering where it came from

OpenStudy (anonymous):

lol, this came from my math professor. I had no idea how to do it either.

OpenStudy (anonymous):

tell your math professor (gently) that you think maybe there is a typo

OpenStudy (anonymous):

it is possible to write \[f(x)=c\] meaning f is constant. answer above is still good. b cannot be 0 or -5

OpenStudy (anonymous):

So the answer would actually be: \[\left\{ x | x \in \mathbb{R} \ except\ x \neq -5\ or\ x \neq 0\right\}\]